OpenStudy (anonymous):

Which statement best describes the effect of replacing the function f(x) = 2x - 2 with the function g(x) = 2x + 5? The graph shifts 7 units left. The graph shifts 5 units left. The graph shifts 2 units up. The graph shifts 3 units up.

OpenStudy (anonymous):

@jagr2713 , @TheSmartOne

OpenStudy (anonymous):

Please help!

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey

OpenStudy (misty1212):

to get from \(2x-2\) to \(2x+5\) you would add 7 that move it up 7 units

OpenStudy (misty1212):

i see that is not one of your choices though, hmm

OpenStudy (anonymous):

Yeah

OpenStudy (misty1212):

let me think, maybe they want you to say it shifts left or right, kinda strange answer 3 and 4 are out for sure

OpenStudy (anonymous):

Yeah I wouldnt think one of those two are it.

OpenStudy (misty1212):

i cannot see another possible answer they both have the same slope, so they are parallel, hard to say how they want you to answer one crosses the x axis at 1, the second crosses the x axis at \(-\frac{5}{2}\) so you could say it was shifted left but not by any of your choices

OpenStudy (misty1212):

ok here is a picture you can see one is 7 units up form the other or \(\frac{7}{2}\) to the left

OpenStudy (anonymous):

Yeah I see that

OpenStudy (misty1212):

i will be happy to help, but i didn't help much on this one maybe it is a typo?

OpenStudy (misty1212):

\[\sqrt2\sqrt[3]{2}\] right? easier to use fractional exponents

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

\[\frac{1}{2}+\frac{1}{3}=\frac{5}{6}\]

OpenStudy (misty1212):

so your answer is \[\sqrt[6]{2^5}\]

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

we have to compute the second year and the fourth year

OpenStudy (misty1212):

\[500(1.05)^4=607.75\] \[500(1.05)^2=551.25\]is a start

OpenStudy (misty1212):

then subtract, and divide the result by 2, since it is the average over two years

OpenStudy (misty1212):

\[\frac{607.75-551.25}{2}\] is what you need

OpenStudy (anonymous):

I got 28.25

OpenStudy (misty1212):

yeah me too

OpenStudy (triciaal):

for the 1st one I would put D 3 units up

OpenStudy (misty1212):

yeah B

OpenStudy (anonymous):

Yay!

OpenStudy (misty1212):

whew !

OpenStudy (anonymous):

How about for this one? I think it is D The sales totals at Linda's food store have increased exponentially over the months. Which of these best shows the sales in the first three months? $1200 in the first month, $1260 in the second month, and $1323 in the third month $1200 in the first month, $1250 in the second month, $1300 in the third month $1200 in the first month, $1272 in the second month, $1344 in the third month $1200 in the first month, $1285 in the second month, $1370 in the third month

OpenStudy (triciaal):